A block weighing 10N is dropped from a height h=12cm above a spring balance with k= 1000N/m. What maximum weight will the balance read momentarily?
1
Expert's answer
2020-05-03T17:16:39-0400
"mg(h+x)=0.5kx^2\to \\10(0.12+x)=0.5(1000)x^2"

"x=0.06\ m"

"F=kx=(1000)(0.06)=60\ N"